A set of charged plates is separated by 2.22*10^-4m. When 5.24*10^-9 C of charge is placed on the plates, it creates a potential difference of 240 V. What is the are of the plates?


1
Expert's answer
2021-05-19T11:03:50-0400

With the provided information, we can find the capacitance of the system:


"C=\frac QV."

On the other hand, the capacitance can be expressed as


"C=\frac{\epsilon_0A}{d},\\\space\\
A=\frac{Cd}{\epsilon_0}=\frac{Qd}{V\epsilon_0}=5.47\text{ cm}^2."